Ops/Comm Scheduling meeting

Location : G06 in R66 at 2pm

Phone access : ReadyTalk number
US 866-740-1260
UK 0800-496-0576
Swiss 0800-70-53-51
Italy 800182592
Access Code: 2155402

  • SS Magnet status
  • Hydrogen status
  • Issues arising from previous running
  • Operations :
    • C&M System Check report (Ajit/Ed)
    • Establishing a timetable for the runup : when should systems (including magnets) be ready for running?
      • Beam schedule
        • Run up on Jan 30 - Feb 12.
        • Start of user cycle : Feb 14
        • Start of data-taking : Feb 15
      • Reserve shifter support from Feb 10 - Feb 14
      • MOM formally starts on Feb 9 (but should be present for the runup activities but will not be on 24-hour call)
      • Possible runup plan (framework for discussion)
        • Monday 6th : Preparation meeting with Ed/Ajit/Colin/Josef/Others to go over readiness for running
        • Tuesday 7th (subject to meeting on Monday) : Bring up magnets to running configuration, take some pulser data and take magnets down
        • Saturday 11th : Expert led run without solenoids + Beam tune
        • Monday 13th : Triage actions arising from expert led run
        • Tuesday 14th : Bring up magnets in the evening and monitor overnight by reserve shift.
        • Wednesday 15th : Data-taking start at 08:00
  • Shift signup for first half of Feb is viable now (but not full)
Shift Block ID # signed up # shifters still needed
1 6 0
2 4 2 Need one overnight shift per night (5 nights) for monitoring
3 4 2 Need one overnight shift per night (5 nights) for monitoring
4 4 1

Magnet monitors do not need to be trained shifters

  • BLOCs are needed : Feb 20 - 24 and Feb 27 - Mar 1
  • AOB:


  • CC Magnet status (J Boehm)
    All 3 CC magnets are now in flip mode; cold and full of helium. Everything is presently being run into shorts for the CC IOC testing.
    Handover from Pierrick is in progress. A few GUIs have been changed.
    Are seeing instability in SSU power supply (into shorts): last week suddenly ramped from 150 to 215 A; lost control of GUI; ramped down manually.
    Similar last night, with control cables disconnected from the AMI controller - again unstable and major changes in current on its own, implies a hardware issue.
    Now running with a different controller - seeing issues still. May possibly be associated with a purely resistive load.
  • Hydrogen status (M Tucker)
    Cooled down in R9 with Neon, made 2 to 2.5 litres of liquid, less than 1 K difference between condenser and absorber.
    Expect about 1 l every 6 hours, so over 5 days to fill with LH2 needing a 24hour rota.
  • Issues arising from previous running (DC)
    Need a comprehensive list of encountered issues, else can't assume they will be addressed before running resumes. We are now well past halfway through the shutdown.
  • Shutdown work in Jan and Feb ( official version )
    Issues/Progress on workplan items (using a stale list):
    • PT100s - done
    • ODH sensor service - done
    • POMPOMs sensor test - done
    • Neutron Monitor: ISIS can't do it; will return to MICE for us to make arrangements. Continue with temporary monitor this cycle.
    • Proton Absorbers - remedial work done
    • "New" chiller plumbing - done
    • #1878 Trench heat exchanger to counter-flow - done, possibly some small leaks
    • Diffuser iris 4: needs next SSU run: Tuesday 7th
      Compressed air: Activair coming to check the outside compressor to see why it didn't come on.
    • #1399 MICEnet lockout - tomorrow morning
    • #1888 Diffuser air - reported done
    • Airflow monitoring changes - done
    • PPS Functional Test - done
    • Earth leakage tests - done
    • New target1ctl - issues with O/S and libraries...
    • R78 - replacement flanges in progress
  • The Decay Solenoid is cold. Only known issue is the episodes of fridge instability seen last run.
    Presently the cooling lead flow is on, to check the stability of the fridge; otherwise ready to power.
  • C&M System Check report (Ajit/Ed)
    CC Magnet UIs/IOCs tested last week - changes have been made on various machines, need to be aggregated.
    Target SM tested today.
    Need to test Run Control: Thursday?
    Need to review the tags listed in the CDB
    The NIKHEF Hall probes have been separated out into a different IOC.
    Most of the hardware stuff is done.
  • Run-up plan
    • Roof water/chiller needs a load test with the air-blast active - tomorrow or Thursday subject to Hall work and state of pipes
    • Monday 6th: Preparation meeting with Ed/Ajit/Colin/Josef/Others to go over readiness for running; remove shorts from link boxes, ATW and other prep.
    • Tuesday 7th: Bring up CC magnets to running configuration, take some pulser data and take magnets down; check Diffuser operation
    • Saturday 11th - Half-day of ISIS beamtime: expert led run without CC Magnets + Beambump tune (Hayley, NB Ed has to leave mid-afternoon)
    • Monday 13th: Triage actions arising from expert led run; BLOC training on everything that isn't target
    • Tuesday 14th: Bring up magnets in the evening and soak-test overnight (reserve shifters).
    • Wednesday 15th: Data-taking start at 08:00
  • Shift signup:
    With the LH2 install going ahead there will be no running during second half of 2016/05 and thus no available shift allocation - MICE that have not yet made shift quota take note.
  • AOB:
    Linde Gas Analyser webcam - where has it gone? It is being PAT tested.


  • JBoehm to get SS operating info from PHanlet (ongoing)
  • JBoehm to get/impose consensus from SC Magnets group on running of "Central CC GUI"
  • DC to collate list of issues arising from previous running
  • Status of miceraid5 disk cleanup (stands)
  • AKurup to make and circulate plan for NIKHEF probe IOC
  • SBoyd - ISIS machine-physics beam time
  • DC to ensure Health Physics have checked temporary neutron monitor (stands)
  • SBoyd - ensure MICE-wide discussion of MOM C&M panel
  • Henry to get gaps in BLOC rota plugged

Updated by Nebrensky, Henry about 7 years ago ยท 21 revisions